# The Relational Origins of Rules in Online Communities

### v1

@@ 6. Discussion
Although online communities often adopt rules for the functional regulation of behavior or content, we find that this is only part of the story. Rules also reinforce a community’s distinctive identity within an online platform and establish a community’s legitimacy in broader community networks. This insight complements challenges prior social computing work from both the “social engineering” and “institutional design” perspectives and reveals underappreciated relational aspects of rulemaking. Some of these relational aspects are external, such as when rules draw distinctions or signal alignments between communities. Others are internal, as when rules reinforce identity or consensus among a community’s leadership team or between leaders and regular members. Both internal and external relationships are important in the processes of rulemaking; they draw attention to problems, help form solutions, and shape decisions about changing the rules—particularly as community leaders have limited capacity and time.
@@ 6.1. Relational Dynamics of Online Community Rules
A prevailing assumption is that online community rules formalize social norms that regulate interactions within communities (Butler et al., 2008; Kraut and Resnick, 2012; Grimmelmann, 2017; Chandrasekharan et al., 2019; Seering, 2019). This assumption treats rules as instruments for directing human interactions in computer-mediated environments. We find that, in addition to advancing this functional purpose, also serve relational purposes as leaders design rules with respect to other communities, leaders, and users.
First, rules set and sustain community boundaries, particularly as a means of protecting community identity amid evolving membership dynamics. The challenge of protecting identity is important—particularly for communities prone to influxes of new members who may alter the community’s composition or undermine its focus (Lin et al., 2017). In response, communities discoverable on large platforms and open to broad participation often implement rules that restrict off-topic content in order to preserve their identity and focus. While prior work has suggested that rules contribute to identity formation (Butler et al., 2008; Fiesler et al., 2018; Seering et al., 2019), our findings show that communities also defend their identity over time by prescribing strict guidelines for participation, often through reference to related communities. We extend prior work by Hwang and Foote (2021) and TeBlunthuis et al. (2022) who emphasize the importance of membership size and niche audiences within ecosystems of interdependent communities. This relational use of rules helps explain the concept of “nested niches,” in which communities thrive through their embeddedness in a broader socio-informational ecosystem (Hwang and Foote, 2021). Rules provide tools for drawing and maintaining boundaries between specialist and generalist spaces—especially on platforms, such as Reddit, where communities often emerge from one another and share overlapping members and topics (Tan, 2018; TeBlunthuis and Hill, 2022).
{mdframed}
@@
Second, rules signal legitimacy to potential newcomers and users—particularly leaders—of other communities. Doing so demonstrates commitment to governance as a means of attracting newcomers and establishing a good community reputation within a platform. This pattern was most salient on the Fediverse where reputation is vital for maintaining connections to other servers in the decentralized network environment. Many Fediverse leaders adopted the same rules as the most popular servers because these were a widely recognized standard for governance, enhancing the credibility of leaders who adopted them among their peers. A community’s rules thus signal to important external audiences that the leadership takes governance (i.e., content moderation) seriously and abides by inter-server social norms.
@@
Such institutional isomorphism is dramatic on the Fediverse, but reflects more general dynamics. Leaders of new online communities, facing uncertainty in how to succeed in their environment (Singh et al., 1986), are prone to failure (Kraut and Resnick, 2012) and may lack requisite resources or knowledge to write well-crafted rules on their own. Instead, they model their community’s institutions (including their rules) on older communities that they perceive to be legitimate or successful (DiMaggio and Powell, 1983). In our study, founding leaders believed that a robust rule set would both attract newcomers and strengthen relations with neighboring communities.
{mdframed}
Design Implication: Community leaders often seek legitimacy by aligning with broader, recognizable norms across the platform. Platforms can facilitate this by providing curated rule templates drawn from similar communities that prompt customization to establish a distinct identity.
Our findings also point to ways that sociotechnical system designs condition rulemaking. Rationales around legitimacy were paramount on the Fediverse but absent from interviews with communities on Fandom and Reddit. Meanwhile, rationales around identity were central on Reddit and Fandom but less salient in interviews with Fediverse leaders. We hypothesize that this difference is attributable to how communities on Reddit and Fandom are strongly held to platform-wide norms and policies, but Fediverse servers depend on connecting to a decentralized network of autonomous communities. Platform-wide norms reduce the uncertainty that creates isomorphic pressures, making the problem of identity differentiation more central. By contrast, because Fediverse communities engage in community-level blocking (Colglazier et al., 2024), it becomes much more important for them to signal legitimacy through their rules and behavior. Doing so indicates whether they are aligned with other communities and allows them to sustain a wide network. No similar affordance for one community to sanction another exists on Fandom and Reddit. As long as Fandom wikis and subreddits adhere to the platform-wide norms, governance signals do not affect their inclusion in the community ecology.
{mdframed}
@@ 6.2. Navigating the Challenges of Rule Change
First, additional problems may trigger a need to change the rules in an online community. Recent quantitative studies of rule changes in subreddits and Minecraft servers have found that increasing membership size and community age associate with leaders deciding to change the rules (Frey and Sumner, 2019; Reddy and Chandrasekharan, 2023). In contrast, we find that periods of decline, such as decreasing activity in the community, may also provoke rule changes. But instead of adding rules to restrict content, leaders may actually remove or relax rules to encourage wider participation.
{mdframed}
@@
Next, we find that leaders, while knowing problems exist, may lack the capacity to implement effective rule changes to resolve those problems. March et al. (2000), analyzing the evolution of written rules at a university, argue that organizations are “thermostatic” in their attention to problems: as long as performance is satisfactory, problems may go unresolved or unnoticed. Online communities, we suggest, may be especially prone to neglect problems because their voluntary leaders have limited resources. Volunteer leaders of online communities do unpaid civic labor to sustain their communities (Dosono and Semaan, 2019; Matias, 2019a; Wohn, 2019) and often quit because of their lack of time brought about psychological distress (Schöpke-Gonzalez et al., 2024). Yet as communities mature, new problems often demand new rules (Butler et al., 2008; Kiene et al., 2016; Lin et al., 2017; Reddy and Chandrasekharan, 2023). Leaders we interviewed cited similar challenges in finding time to change their rules, carefully weighed the demands of potential new rule enforcement, and in one case, replaced the leadership team with members better positioned to improve community governance.
{mdframed}
@@
One way to reduce strain on leaders is to share governance work with the broader community. Doing so also taps into community members’ collective knowledge to help in identifying problems and co-creating solutions (Turco, 2016; Hwang et al., 2024). Through Frey and Schneider’s (Frey and Schneider, 2021) and Schneider’s (Schneider, 2022) participatory governance lens, none of the communities we investigated for this study encoded mechanisms for direct, democratic participation in governance; none had binding mechanisms through which community members’ voice directly influenced decision-making by leaders. Fediverse leaders we interviewed were particularly autocratic: they rarely changed their rules, never held discussions or encouraged feedback about the rules and governance, and made decisions with no input from members. In fact, Fediverse leaders often said they wanted to create their own server where they would get to decide how things should be. In contrast, Fandom leaders expressed concerns about managing delicate relationships with high quality wiki contributors and gathered their input before making decisions. Subreddit leaders seemed even more involved with and accountable to their members. Many reported rulemaking processes that involved much more participation from members and even “soft” binding mechanisms for member demands.
What makes Reddit so different? We propose that a combination of sociotechnical affordances and institutionalized norms sustain a stronger culture of participatory governance. On the affordances side, centralized discussion spaces and integrated voting tools help members collectively communicate their concerns and values to leaders, increasing pressure on them to solve problems. On the normative side, Reddit users have developed sitewide values emphasizing individual freedom and a deep skepticism of moderators who act unaccountably. While nothing binds subreddit moderators to accept member demands, soft binding mechanisms, along with low costs of exit, constrain moderators. For example, several subreddit leaders we interviewed were cautious about making decisions without input from their community members. They feared that if influential members’ concerns weren’t taken seriously, these members would take their expertise elsewhere. Similar examples with Fediverse leaders were non-existent in our sample.
Taken together, this arrangement may result in leader-member relationships that are mutually dependent: members need leaders to moderate content and enforce rules, while leaders need members to contribute content and sustain community activity levels. Evoking metaphors in content moderation from Seering et al. (2020), the leaders of these subreddits may be more like “governors” who satisfy the needs of their constituents and less like the “custodian” wiki leaders who curate and tidy wikis, or the Fediverse “dictators” who create and run servers however they see fit.
@@ 6.3. Limitations and Future Work
As with all interview-based studies, our findings are shaped by the scope of our sample, participants’ subjectivity and self-presentation, and the interpretative nature of our qualitative analysis. While our sampling strategy aimed to capture a diverse and meaningful range of communities across three different sociotechnical systems, our findings are not intended to be representative or generalizable. Instead, our findings instead provide in-depth insights into rulemaking practices among online community leaders. We only looked at Reddit, Fandom, and the Fediverse when we could have considered other platforms as empirical settings. On the other hand, very few platforms besides those three have publicly available methods of constructing datasets of community rules. Future work could extend or contribute to our findings by interviewing leaders from communities on other platforms. Because we exclusively interviewed leaders (i.e., admins and moderators), future work is needed to understand community members’ perspectives and experiences related to rules, especially with respect to how they select communities to join, their perceptions of legitimate community governance, and how they navigate the challenges of voicing problems in online communities to inspire change. We proposed that institutional isomorphism may be a mechanism guiding the behaviors of leaders to adopt similar rules, but longitudinal studies of rule diffusion across community networks should evaluate such a claim. Additionally, an ecological analysis of community rules could investigate whether and how rules are used to construct niches within platforms or more complex community ecosystems. Although recent work is beginning to analyze rule changes in online communities, we also propose further statistical work is needed to investigate the relationship between rule changes and member participation.
@@ Appendix A Rules in subreddits, Fandom wikis, and Fediverse servers
This section provides examples of rules from the communities from which we interviewed leaders. This first set of rules is from the subreddits r/cybersecurity (Figure 1). The second set of rules come from the Marvel Fandom wiki (Figure 2). Finally, the third example of rules comes from the SFBA.social Fediverse community (Figure 3). Communities on all three platforms tended to list additional detail about their rules, usually on other pages.
Figure 1. Rules for the r/cybersecurity subreddit community.
Figure 2. Rules for the Marvel Fandom wiki community.
Figure 3. Rules for the SFBA.social Fediverse community.
@@ Appendix B Participating Community Characteristics
This section gives basic descriptions of the online communities we selected to sample interview participants. We give the community’s name, when it was created, an estimate of its member population size, and the number of rules the community had adopted at the time of data collection.
Community Created Size Rules
r/kdramarecommends 2014 68889 11
r/explainlikeimfive 2011 20945799 10
r/Supernatural 2008 202207 9
r/cybersecurity 2012 381366 9
r/huntsvillealabama 2008 60028 8
r/diabetes_t1 2013 41638 8
r/amcstocks 2021 67428 15
r/harrypotter 2008 1324162 14
r/takeaplantleaveaplant 2011 51156 15
@@
Community Created Size Rules
Aesthetics Fandom 2018 4652 18
Marvel Fandom 2005 23481 41
Naruto Fanon Fandom 2007 4588 59
Valorant 2020 880 21
DC Wiki 2009 15025 33
Riordan Wiki 2007 13365 135
Wings of Fire Fanon 2013 10125 45
Brickpedia 2006 11926 28
@@
Community Created Size Rules Software
sfba.social 2019 3700 7 Mastodon
toot.wales 2018 981 3 Mastodon
medievalist.masto.host 2018 26 4 Mastodon
freeradical.zone 2017 278 15 Mastodon
bostonmusic.online 2020 2 6 Mastodon
neue.city Not Available 60 22 Pleroma (Akkoma)
hackers.town 2017 211 16 Mastodon
ausglam.space 2018 235 9 Mastodon
vocalounge.cafe 2020 49 6 Mastodon
mspsocial.net 2018 370 11 Mastodon
sonomu.club 2019 79 10 Mastodon
mastodon.radio 2018 1100 9 Mastodon
expired.mentality.rip Not Available Not Available Not Available Pleroma
vantil.refchat.net Not Available 31 Not Available Pleroma
mk.disqordia.space Not Available Not Available Not Available Misskey
glammr.us 2018 271 3 Mastodon

### v2

@@ 6. Discussion
Although online communities often adopt rules for the functional regulation of behavior or content, we find that this is only part of the story. Rules also reinforce a community’s distinctive identity within an online platform and establish a community’s legitimacy in broader networks. This insight complements and challenges prior social computing work from both the “social engineering” and “institutional design” perspectives and reveals underappreciated relational aspects of rulemaking. Some of these relational aspects are external, such as when rules draw distinctions or signal alignments between communities. Others are internal, as when rules reinforce identity or consensus among a community’s leadership team or between leaders and regular members. Both internal and external relationships are important in the processes of rulemaking; they draw attention to problems, help form solutions, and shape decisions about changing the rules.
@@ 6.1. Relational Dynamics of Online Community Rules
A prevailing assumption in prior scholarship is that online community rules formalize social norms that regulate interactions within communities (Butler et al., 2008; Kraut et al., 2012; Grimmelmann, 2017; Chandrasekharan et al., 2019; Seering, 2019). This assumption treats rules as instruments for directing human interactions in computer-mediated environments. We find that, in addition to advancing this functional purpose, rules also serve relational purposes as leaders design rules with respect to other communities, leaders, and users. We unpack this finding into four design implications.
First, rules set and sustain community boundaries, particularly as a means of protecting community identity amid evolving membership dynamics. The challenge of protecting identity is important—particularly for communities prone to influxes of new members who may alter the community’s composition or undermine its focus (Lin et al., 2017). In response, communities discoverable on large platforms and open to broad participation often implement rules that restrict off-topic content. While prior work has suggested that rules contribute to identity formation (Butler et al., 2008; Fiesler et al., 2018; Seering et al., 2019), our findings show that communities also defend their identity over time by prescribing strict guidelines for participation, often through reference to related communities. We extend prior work by Hwang and Foote (2021) and TeBlunthuis et al. (2022) who emphasize the importance of membership size and niche audiences within ecosystems of interdependent communities. This relational use of rules helps explain the concept of “nested niches,” in which communities thrive through their embeddedness in a broader socio-informational ecosystem (Hwang and Foote, 2021). Rules provide tools for drawing and maintaining boundaries between specialist and generalist spaces—especially on platforms, such as Reddit, where communities often emerge from one another and share overlapping members and topics (Tan, 2018; TeBlunthuis and Hill, 2022; TeBlunthuis, 2025).
@@
Second, rules signal legitimacy to potential newcomers and users—particularly leaders—of other communities. Doing so demonstrates commitment to governance as a means of attracting newcomers and establishing a good community reputation within a platform. This pattern was most salient on the Fediverse where reputation is vital for maintaining connections to other servers. Many Fediverse leaders adopted the same rules as the most popular servers because these were a widely recognized standard for governance. A community’s rules thus signal to important external audiences that the leadership takes governance (i.e., content moderation) seriously and abides by shared social norms.
@@
Such institutional isomorphism is dramatic on the Fediverse, but reflects more general dynamics. Leaders of new online communities, facing uncertainty in how to succeed in their environment (Singh et al., 1986), are prone to failure (Kraut et al., 2012) and may lack requisite resources or knowledge to write well-crafted rules on their own. They might model their community’s institutions (including their rules) on communities that they perceive to be legitimate or successful (DiMaggio and Powell, 1983). In our study, founding leaders believed that a robust rule set would both attract newcomers and strengthen relations with neighboring communities.
Design Implication: Community leaders often seek legitimacy by aligning with broader, recognizable norms across the platform. Platforms can facilitate this by providing curated rule templates drawn from similar communities.
Our findings also point to ways that sociotechnical system designs condition rulemaking. Rationales around legitimacy were paramount on the Fediverse but absent from interviews with communities on Fandom and Reddit. Meanwhile, rationales around identity were central on Reddit and Fandom but less salient in interviews with Fediverse leaders. We hypothesize that this difference is attributable to how communities on Reddit and Fandom are strongly held to platform-wide norms and policies, while Fediverse servers depend on connecting to a decentralized network of autonomous communities. Platform-wide norms reduce the uncertainty that creates isomorphic pressures, making the problem of identity differentiation more central. By contrast, because Fediverse communities engage in community-level blocking (Colglazier et al., 2024), it becomes much more important for them to signal legitimacy through their rules and behavior. No similar affordance for one community to sanction another exists on Fandom and Reddit.
@@ 6.2. Navigating the Challenges of Rule Change
First, additional problems may trigger a need to change the rules in an online community. Recent quantitative studies of rule changes in subreddits and Minecraft servers have found that increasing membership size and community age associate with leaders deciding to change the rules (Frey and Sumner, 2019; Reddy and Chandrasekharan, 2023). In contrast, we find that periods of decline, such as decreasing activity in the community, may also provoke rule changes. However, instead of adding rules to restrict content, leaders may actually remove or relax rules to encourage wider participation.
@@
Next, we find that leaders, while knowing problems exist, may lack the capacity to implement effective rule changes to resolve those problems. March et al. (2000), analyzing the evolution of written rules at a university, argue that organizations are “thermostatic” in their attention to problems: as long as performance is satisfactory, problems may go unresolved or unnoticed. Self-governed online communities may be especially prone to neglect problems because their leaders are volunteers. These leaders are typically unpaid (Dosono and Semaan, 2019; Matias, 2019a; Wohn, 2019) and frequently quit because of their lack of time brings about psychological distress (Schöpke-Gonzalez et al., 2024). Yet as communities mature, new problems often demand new rules (Butler et al., 2008; Kiene et al., 2016; Lin et al., 2017; Reddy and Chandrasekharan, 2023). Leaders we interviewed cited similar challenges in finding time to change their rules, carefully weighed the demands of potential new rule enforcement, and in one case, replaced the leadership team with members better positioned to improve community governance.
@@
One way to reduce strain on leaders is to share governance work with the broader community. Doing so also taps into community members’ collective knowledge to help in identifying problems and co-creating solutions (Turco, 2016; Hwang et al., 2024). None of the communities we investigated for this study were fully democratic in the sense that none had binding mechanisms through which community members’ voice directly influenced decision-making by leaders (Frey and Schneider, 2021; Schneider, 2022). Fediverse leaders we interviewed were particularly autocratic: they rarely changed their rules, never held discussions or encouraged feedback about the rules and governance, and made decisions with no input from members. In fact, Fediverse leaders often said they wanted to create their own server where they would get to decide how things should be. In contrast, Fandom leaders expressed concerns about managing delicate relationships with high quality wiki contributors and gathered their input before making decisions. Subreddit leaders seemed even more involved with and accountable to their members. Many reported rulemaking processes that involved much more participation from members.
What makes Reddit so different? We propose that a combination of sociotechnical affordances and institutionalized norms sustain a stronger culture of participatory governance. On the affordances side, centralized discussion spaces and integrated voting tools help members collectively communicate their concerns and values to leaders, increasing pressure on them to solve problems. On the normative side, Reddit users have developed sitewide values emphasizing individual freedom and a skepticism of moderators who act unaccountably. While nothing binds subreddit moderators to accept member demands, soft binding mechanisms, along with low costs of exit, constrain moderators. For example, several subreddit leaders we interviewed were cautious about making decisions without input from their community members. They feared that if influential members’ concerns weren’t taken seriously, these members would take their expertise elsewhere. Similar examples with Fediverse leaders were non-existent in our sample.
Taken together, these arrangements in Reddit may result in mutually dependent leader-member relationships: members need leaders to moderate content and enforce rules, while leaders need members to contribute content and sustain community activity. Evoking metaphors in content moderation from Seering et al. (2020), the leaders of these subreddits may be more like “governors” who satisfy the needs of their constituents and less like the “custodian” wiki leaders who curate and tidy wikis, or the Fediverse “dictators” who create and run servers however they see fit.
@@ 6.3. Limitations and Future Work
As with all interview-based studies, our findings are shaped by the scope of our sample, participants’ subjectivity and self-presentation, and the interpretative nature of our qualitative analysis. While our sampling strategy aimed to capture a diverse and meaningful range of communities across three different sociotechnical systems, we cannot claim that our findings are representative or generalizable. Instead, our findings instead provide in-depth insights into rulemaking practices of a range of online community leaders. We only looked at Reddit, Fandom, and the Fediverse when we could have considered other platforms as empirical settings. On the other hand, few platforms besides those three have publicly available methods of constructing datasets of community rules. Future work could extend or contribute to our findings by interviewing leaders from communities on other platforms. In 2022, changes to Twitter/X corresponded with a migration of many users to the Fediverse, which may have affected rule adoption patterns and broader norms in governance in that setting while we were interviewing leaders.
Because we exclusively interviewed leaders (i.e., admins and moderators), future work is needed to understand community members’ perspectives and experiences related to rules, especially with respect to how they select communities to join, their perceptions of legitimate community governance, and how they navigate the challenges of voicing problems in online communities to inspire change. We proposed that institutional isomorphism may be a mechanism guiding the behaviors of leaders to adopt similar rules, but longitudinal studies of rule diffusion across community networks should evaluate such a claim. Additionally, an ecological analysis of community rules could investigate whether and how rules are used to construct niches within platforms or more complex community ecosystems. Although recent quantitative work has analyzed rule changes in online communities, we also propose that further statistical work is needed to investigate the relationship between rule changes and member participation.
@@ Appendix A Rules in subreddits, Fandom wikis, and Fediverse servers
This section provides examples of rules from online communities in each of the three settings of our study. This first set of rules is from a subreddit community (Figure 1). The second set of rules come from a Fandom wiki community(Figure 2). Finally, the third example of rules comes from a Fediverse community (Figure 3). Communities on all three platforms tended to list additional detail about their rules, usually on other pages.
Figure 1. Rules for a subreddit community.
Figure 2. Rules for a Fandom wiki community.
Figure 3. Rules for a Fediverse community.
@@ Appendix B Participating Community Characteristics
This section gives basic descriptions of the online communities we selected to sample interview participants. We give the community’s topic, when it was created, an estimate of its member population size, and the number of rules the community had adopted at the time of data collection.
Community Topic Created Size Rules
Television genre 2014 68889 11
Q&A 2011 20945799 10
Television series 2008 202207 9
IT professionals 2012 381366 9
City-specific discussions 2008 60028 8
Illness support 2013 41638 8
Finance 2021 67428 15
Book and film series 2008 1324162 14
Hobbyist exchanges 2011 51156 15
@@
Community Topic Created Size Rules
Fashion Styles 2018 4652 18
Comic book franchise 2005 23481 41
Anime fanfiction lore 2007 4588 59
Multiplayer video game 2020 880 21
Comic book franchise 2009 15025 33
Book series 2007 13365 135
Book series 2013 10125 45
Toy franchise 2006 11926 28
@@
Community Topic Created Size Rules Software
City-specific and general discussions 2019 3700 7 Mastodon
Region-specific and general discussions 2018 981 3 Mastodon
History 2018 26 4 Mastodon
General discussions and news 2017 278 15 Mastodon
City-specific and music discussions 2020 2 6 Mastodon
General discussions and news Not Available 60 22 Pleroma (Akkoma)
Computer hacking 2017 211 16 Mastodon
Region-specific and museum discussions 2018 235 9 Mastodon
Music equipment discussions 2020 49 6 Mastodon
City-specific and general discussions 2018 370 11 Mastodon
Musicians 2019 79 10 Mastodon
Amateur hobbyists 2018 1100 9 Mastodon
Video sharing Not Available Not Available Not Available Pleroma
Religion-specific discussions Not Available 31 Not Available Pleroma
General discussions and news Not Available Not Available Not Available Misskey
Museum discussions 2018 271 3 Mastodon